# Extraction of Singular Patterns from a Vector Field via Persistent Path Homology

### v2

@@ IV-B Extracting Position of Singularity
Here, $(x_{i},y_{i})$ represents the midpoint coordinates of the four edges of the square, with corresponding weights $w_{i}$ for $i=1,2,3,4$. The total weight is denoted as $w$, calculated as $w=\sum_{i=1}^{4}w_{i}$. When considering two parallel edges within the square, the weight center is closer to the edge with a greater weight, similar to the singularity. Consequently, horizontally and vertically, the weight center and the singularity are positioned on the same side of the geometric center of the square. Therefore, if we use the notation $\varepsilon$ to denote the length of the square’s edges, the errors in the $x$ and $y$ coordinates between the weight center and the singularity are less than $\varepsilon/2$. The procedure for extracting the position of singularity is included in Algorithm 1 illustrated in subsection IV-D.
IV-C Segmenting the Region of Singular Pattern
Another crucial issue is the segmentation of a region exhibiting a singular pattern. For instance, when tracking a tropical cyclone, it is desirable to determine not only its central location but also the area it will impact. In our study, we utilize singular polygons as an approximation for delineating a region of the singular pattern. This concept is inspired by the methods used to extract representative cycles of the homology groups from classical persistent homology theory [36].
Let us revisit the definition of an angle-based digraph, where the edge weights in $G$ represent absolute value of rotation angles. Considering the characteristics of singular patterns, we observe that the rotation angle between vectors at the ends of an edge diminishes as we move farther away from the singularity. Consequently, within the digraph filtration, in the vicinity of the singularity, the edge that is more distant from the singularity possesses a smaller weight and be added earlier than the closer edge.
To streamline our discussion, we introduce the concept of a reduced digraph, which facilitates the examination of path homology within a digraph.
Definition 12.
(Isolated edge) If an edge in graph $G$ contains a vertex with degree 1, it is an isolated edge in graph $G$.
Definition 13.
(Reduced digraph) Let $G$ be a digraph, $G_{r}$ is a subgraph of $G$, then we say $G_{r}$ is a reduced digraph of $G$ if $G_{r}$ is the maximal subgraph of $G$ that does not contain any isolated edges.
Theorem 7.
[3] Suppose a graph $G$ has a vertex $a$ such that there is only one outcoming edge $a\rightarrow b$ from $a$ and no incoming edges to $a$. Let $G^{\prime}=\left(V^{\prime},E^{\prime}\right)$ be the digraph with $V^{\prime}=V\backslash\{a\}$ and $E^{\prime}=E\backslash\{ab\}$. Then $H_{p}(G)\cong H_{p}\left(G^{\prime}\right)$ for all $p\geq 0$. The theorem still holds if we change the condition of edge to $b\rightarrow a$.
Figure 6: The digraph on the right is the reduced digraph of the one on the left.
Figure 7: (a) Example of a discrete vector field that has a singular pattern. (b) four selected moments in the reduced digraph filtration, where the pink region in the second one is the singular polygon corresponding to the singularity in the vector field on the left.
According to Theorem 7, a digraph and its reduced digraph have the same path homology group. For instance, in Fig. 6, the digraph on the right is the reduced digraph of the one on the left. Their path homology groups are identical. Therefore, if our focus is solely on the path homology of a digraph, we can equivalently study its reduced digraph.
Based on the preceding discussion, when adding the edges in a non-decreasing order of weight, there exists an edge that forms the minimum polygon encompassing the singularity we are seeking. In other words, after adding an edge $e$ with weight $w(e)$, a new path homology generator is created, which can be represented by the minimum polygon containing $e$.
An example is shown in Fig. 7. On the left, a discrete vector field is depicted, showing a singular pattern generated by a singularity. On the right, four moments are selected from the reduced digraph filtration. The pink region in the second moment represents the minimum polygon encompassing the singularity, and the minimum polygon is situated within the region where the original vector field exhibits the singular pattern.
Definition 14 (Singular Polygon).
The minimum polygon described above is referred to as the singular polygon corresponding to the singularity in $X$.
The singular polygon serves as a neighborhood of the singularity, illustrating the singular pattern.
We now discuss the process of obtaining the singular polygon using persistent path homology. When the aforementioned edge $e$ is added, a new point $(w(e),+\infty)$ appears in $Dgm_{1}$. Therefore, we can identify this edge by examining all the points in $Dgm_{1}$ in the form of $(a_{i},+\infty)$. By traversing these points in non-decreasing order of the value of $a_{i}$, we can locate the edge $e_{i}$ in $G$ with weight $a_{i}$ and analyze the reduced digraph $G^{a_{i}}_{r}$. If $w(e)$ represents the minimum value at which the singular polygon appears in $G^{w(e)}_{r}$, then the corresponding edge $e$ is responsible for the birth of this singular polygon. Once this edge $e$ is identified, we record the singular polygon and draw the reduced digraph $G^{w(e)}_{r}$ to give a visualization of it. This process is also included in Algorithm 1.
IV-D Algorithm and Complexity Analysis
Now we present our algorithm for extracting singularities and singular polygons from a planar discrete vector field.
Algorithm 1 Extracting Singularity and Singular Polygon
@@ IV-D Algorithm and Complexity Analysis
5:  For each point $(w,+\infty)$ in $I$ do    If $G^{w}_{r}$ represents an earliest state in reduced digraph filtration forming a minimal polygon $P$ with index 1 then       Record the polygon $P$.    End ifEnd for
5:  The singularities and the corresponding singular polygons.
In analyzing the complexity of the Algorithm 1, we assume the input $X$ has a rectangular grid of points with $m$ points in each row and $n$ points in each column (hence $mn$ grid points in $X$ in total), and in $Dgm_{1}(G)$ there are $k$ points persist to +$\infty$+ ∞. Note that $k$ represents the number of non-boundary squares in $G$, thus is not greater than $mn$. We consider the time complexity of each steps. Step 1 has a complexity of $O(mn)$, which represents a linear relationship between the input sizes $m$ and $n$. The computation persistent path homology of planar digraph in step 2 exhibits a complexity of $O((mn)^{\omega})$ [11], where $\omega$ denotes the exponent of matrix multiplication, currently known to be less than $2.373$. This step is particularly significant due to the high exponent, which can dominate the overall complexity for large input sizes. The sort process in step 3 has a complexity of $O(k\log k)$. Step 4 has a complexity of $O(kmn)$, but it does not exceed $O((mn)^{\omega})$ since $k$ is controlled by $mn$ and is much smaller than $mn$ in general cases. Lastly, in step 5, the complexity of searching the polygon $P$ and computing its index is also $O(kmn)$, as it may repeat at most $k$ times, and each search may involve traversing the entire grid to identify the polygon. Considering the relative magnitudes of these complexities, the step of computing persistent path homology with $O((mn)^{\omega})$ stands out as the dominant term, determining the overall complexity of the algorithm. Therefore, if there are $N$ grid points in $X$, the overall time complexity of Algorithm 1 is $O((N)^{\omega})$.
V Differences Measurement of Vector Fields
@@
VI Experiments and discussions
All the experiments in this section were conducted on a server equipped with a 2.90GHz Intel Core i7-10700 CPU and 16.0 GB RAM.
We will first present the experimental results obtained by using our method to track two tropical cyclones, namely Khanun and Saola, which emerged in 2023 and exhibited unusual trajectories. In this experiment, we use wind field datasets obtained from the Remote Sensing Systems Cross-Calibrated Multi-Platform’s 6-hourly ocean vector wind analysis product on a 0.25 deg grid [37]; it is a gridded Level 4 (L4) product that provides vector wind over the world’s oceans and provides gridded wind vector data at 4 times of day: 00:00Z, 06:00Z, 12:00Z, and 18:00Z (here Z stands for Universal Time Coordinated, UTC).
Various techniques can be used to detect the center of tropical cyclones [39]. For example, mathematical morphology operations applied to satellite cloud images [40, 41, 42, 43], methods based on vector field [44, 45, 46] or scalar field [47], and AI-based methods utilizing satellite cloud images [48, 49]. These methods may involve multiple inputs and techniques, such as various dynamic and thermodynamic variables at different altitudes, and can be complex. In contrast, our method only utilizes wind field data to analyze rotational patterns and determine the center positions of cyclones, highlighting one of our advantages. Moreover, our method can also measure topological changes in tropical cyclones over time, demonstrating an additional advantage of the approach. Another experiment utilized the geomagnetic field datasets obtained from the International Geomagnetic Reference Field (IGRF) [38], a mathematical model describing the geomagnetic field and its changes, to determine the position of Earth’s dip poles. Additionally, the algorithm proposed in [11] is used for computing one-dimensional persistent path homology.
VI-A Experiment Results of Extracting Singular Patterns
Figure 8: Pipeline of our procedure. An illustrative example is presented, where we use the satellite cloud image of typhoon Khanun to display the wind field data of the area that is selected in our experiments.
@@ VI-A Experiment Results of Extracting Singular Patterns
Figure 9: Center tracking results of two typhoons. (a) Results of typhoon Khanun from 2023/07/31 to 2023/08/08. (b) Results of typhoon Saola from 2023/08/25 to 2023/09/01.
@@
In the first experiment, we determine the centers of tropical cyclones based on the wind fields. Fig. 8 shows the pipeline of our procedure. As an illustrative example, we chose a specific local area from the wind field of typhoon Khanun. By processing the wind field data, we acquire the approximate latitude and longitude of the center positions of tropical cyclone for each provided moment. These positions are then plotted on a single map, as depicted in Fig. 9 (a) displays the positioning result for Khanun, whereas Fig. 9 (b) presents the positioning result for Saola. Each point on the map is labeled with the corresponding date, and the color indicates the observation time. The obtained positioning longitudes and latitudes for the centers of Khanun and Saola are presented in Table IV, alongside the actual longitudes and latitudes, as well as the errors. Here the actual positions of the tropical cyclones are obtained from the Regional Specialized Meteorological Center for Marine Meteorological Service, Beijing [^1] http://eng.nmc.cn/typhoon/ .
In addition, we compute the singular polygon that encompasses the center of a tropical cyclone and depict it in the corresponding reduced digraph. The wind field data from 2023-08-02-00:00 UTC is selected, covering a longitude range of $123.625^{\circ}$–$133.625^{\circ}$ and a latitude range of $21.125^{\circ}$–$31.125^{\circ}$. The pipeline of our process is illustrated in Fig. 8. By executing Algorithm 1, we generate the digraph displayed in Fig. 10. In this figure, the singular polygon containing the center is highlighted in pink. Based on our previous discussion, this polygon should represent a region where the singular pattern exists, indicating the area influenced by the tropical cyclone Khanun. Considering that each pair of adjacent vertices is 0.25 degrees apart, we can determine, using the latitude-longitude to kilometer conversion formula [^2] For latitude, 1 deg $\approx$ 110.574 km; for longitude, 1 deg 111.320$\approx$ 111.320$\times$ 111.320 × cos(latitude) km, that $P$ falls within the coverage of the tropical cyclone. This determination is supported by the data of wind circle radius shown in Table I, obtained from the Regional Specialized Meteorological Center for Marine Meteorological Service in Beijing. Thus, we confirm that the singular polygon represents a region where the singular pattern exists. However, given noise in the field, the shape of the singular polygon may exhibit some irregularity.
TABLE I: Four Direction Wind Circle Radius of Khanun at 2023-08-02-00:00 UTC
NE SE SW NW
[…]
Figure 10: Segmentation of singular polygon using the wind field data at 2023-08-02-00:00 UTC. The pink area shows the singular polygon that corresponds to the tropical cyclone center, which is an area influenced by the tropical cyclone.
TABLE II: Positioning Results of Magnetic Dip Poles and Errors

### v3

@@ 4.2 Determining the position of singularities
Here, $(x_{i},y_{i})$ represents the midpoint coordinates of the four edges of the square, with corresponding weights $w_{i}$ for $i=1,2,3,4$. The total weight is denoted as $w$, calculated as $w=\sum_{i=1}^{4}w_{i}$. Hence, the center can only be in the diamond formed by the edge midpoints of the square’s edges. When considering two parallel edges within the square, the weight center is closer to the edge with a greater weight, similar to the singularity. Consequently, horizontally and vertically, the weight center and the singularity are positioned on the same side of the geometric center of the square. Therefore, if we use the notation $\varepsilon$ to denote the length of the square’s edges, the errors in the $x$ and $y$ coordinates between the weight center and the singularity are less than $\varepsilon/2$. The procedure for extracting the position of singularity is included in Algorithm 1, illustrated in subsection 4.3.
4.3 Algorithm and complexity analysis
Our algorithm for extracting singularities from a planar discrete vector field can be described as the Algorithm 1. To analyze its complexity, we assume the input $X$ has a rectangular grid of points with $m$ points in each row and $n$ points in each column (hence $mn$ grid points in $X$ in total), and in $Dgm_{1}(G)$ there are $k$ points persist to +$\infty$+ ∞. Note that $k$ represents the number of non-boundary squares in $G$, thus is not greater than $mn$. We consider the time complexity of each steps. Step 1 has a complexity of $O(mn)$, which represents a linear relationship between the input sizes $m$ and $n$. The computation persistent path homology of planar digraph in step 2 exhibits a complexity of $O((mn)^{\omega})$ [11], where $\omega$ denotes the exponent of matrix multiplication, currently known to be less than $2.373$. This step is particularly significant due to the high exponent, which can dominate the overall complexity for large input sizes. The sort process in step 3 has a complexity of $O(k\log k)$. Step 4 has a complexity of $O(kmn)$, but it does not exceed $O((mn)^{\omega})$ since $k$ is controlled by $mn$ and is much smaller than $mn$ in general cases. Considering the relative magnitudes of these complexities, the step of computing persistent path homology with $O((mn)^{\omega})$ stands out as the dominant term, determining the overall complexity of the algorithm. Therefore, if there are $N$ grid points in $X$, the overall time complexity of Algorithm 1 is $O((N)^{\omega})$.
Algorithm 1 Extracting Singularities of a Vector Field
@@ 4.3 Algorithm and complexity analysis
4:  The singularities of the vector field $X$.
5 Differences Measurement of Vector Fields
@@
Fig. 5: Pipeline of our method for analyzing the singular patterns of vector fields.
Fig. 6: The positioning results of north dip pole (left) and south dip pole (right) for the years 2000, 2005, 2010, 2015, and 2020.
@@
Fig. 7: Center tracking results of two typhoons. (a) Results of typhoon Khanun from 2023/07/31 to 2023/08/08. (b) Results of typhoon Saola from 2023/08/25 to 2023/09/01.
6 Applications and discussions
All the experiments in this section were conducted on a server equipped with a 2.90GHz Intel Core i7-10700 CPU and 16.0 GB RAM. Additionally, Fig. 5 shows the pipeline of our procedure, and the algorithm proposed in [11] is used for computing one-dimensional persistent path homology.
6.1 Positioning magnetic dip poles
To begin with, we present the application of our method to determine the position of Earth’s magnetic dip poles using the geomagnetic field datasets obtained from the International Geomagnetic Reference Field (IGRF) [33], a mathematical model describing the geomagnetic field and its changes. The magnetic dip poles are defined as the locations where the main magnetic field as a whole is normal to Earth’s surface, which provides a crucial reference for local orientation when navigating on or close to Earth’s surface at high-latitudes [33]. However, observations have shown that the positions of the two magnetic dip poles change from year to year [33], so it is necessary to calculate and update their positions in time. Our approach is well-suited to address this issue. When we project geomagnetic field data onto a map, resulting in a planar vector field, the two dip poles (north dip pole and south dip pole) become singularities of this planar vector field, since the projected vectors are zero at these points. In this experiment, we select two domains around the two dip poles respectively and choose geomagnetic field data at zero altitude. After projecting onto the map, discrete vector fields sampled around the two dip poles are obtained. Using our method, we determine the locations of the two dip poles for the years 2000, 2005, 2010, 2015, and 2020, and compare our results with the data of the actual location provided by IGRF [33]. These results are available in the Appendix B (Table. 2). It is evident that our method incurs only minor positioning errors, thus reaffirming its effectiveness. The positioning results are also plotted on maps (Fig. 6).
6.2 Tracking the center of typhoons
@@
Another application is tracking the center of two tropical cyclones, namely Khanun and Saola, which emerged in 2023 and exhibited unusual trajectories. Here we use wind field datasets obtained from the Remote Sensing Systems Cross-Calibrated Multi-Platform’s 6-hourly ocean vector wind analysis product on a 0.25 deg grid [34]; it is a gridded Level 4 (L4) product that provides vector wind over the world’s oceans and provides gridded wind vector data at 4 times of day: 00:00Z, 06:00Z, 12:00Z, and 18:00Z (here Z stands for Universal Time Coordinated, UTC). Based on these wind fields, we determine the centers of tropical cyclones (which can be regarded as spiral sinks). By processing the wind field data, we acquire the approximate latitude and longitude of the center positions of the tropical cyclone for each provided moment. These positions are then plotted on a single map, as depicted in Fig. 7 (a) displays the positioning result for Khanun, whereas Fig. 7 (b) presents the positioning result for Saola. Each point on the map is labeled with the corresponding date, and the color indicates the observation time. The obtained positioning longitudes and latitudes for the centers of Saola and Khanun are provided in Appendix B (Table. 3 and Table. 4), alongside the actual longitudes and latitudes, as well as the errors. Here, the actual positions of the tropical cyclones are obtained from the Regional Specialized Meteorological Center for Marine Meteorological Service, Beijing [^1] http://eng.nmc.cn/typhoon/.
6.3 Vector field difference measurement
In the following, we will use the wind field datasets that were used in the previous subsection as an example to demonstrate how our method can measure the variation of time-varying vector fields. Similar to the work [35], we will use the resulting persistence diagrams obtained by wind fields to measure the variations of singular patterns of the wind fields of the two tropical cyclones, respectively.
From the blue line in Fig. 8, it is evident that the bottleneck distance reached a local minimum at the moment 2023-08-03-18:00 UTC, coinciding with the time when Khanun became stationary near a position since a notable change in trajectory was taking place. This period exhibited reduced variability in the wind field, resulting in a smaller difference between the persistence diagrams. Similarly, in the orange line in Fig. 8, at the moment 2023-08-27-18:00 UTC, Saola became stationary near a position where its trajectory was changing direction, indicating small variability between the persistence diagrams of the corresponding wind fields. Similar features of bottleneck distance can be observed at other moments, such as 2023-08-31-06:00 UTC, during which period Saola moves slowly within the vicinity of a location. Conversely, if a tropical cyclone undergoes significant movement or multiple turns over a period of time, the corresponding bottleneck distance will be larger. Moreover, a significant fluctuation of the bottleneck distance may indicate that the tropical cyclone is undergoing unstable movement and change.
Fig. 8: Bottleneck distances computed by one-dimensional persistence diagrams derived from Khanun and Saola datasets respectively.
Fig. 9: Robustness of our method to noise. (a) A vector field is presented by its streamlines, and six center points (red) and two saddles (blue) are successfully extracted by our method. (b) Noise is added in the vector field in (a), while our method can keep identifying the singularities successfully. (c) Comparison of singularities identified by triangular linear interpolation method [23] from (b). It fails to detect some singularities while incorrectly identifying a spurious singularity.
Fig. 10: The streamlines of a vector field with 10,000 vectors. (a) Singularities identified by our method. The detected singularities with positive index are drawn in red, while singularities with negative index are drawn in blue. (b) Comparison of singularities identified by triangular linear interpolation method [23]. It fails to detect some singularities while incorrectly identifying some spurious singularities.
6.4 Robustness and comparisons
In this section, we first demonstrate the robustness of our method for detecting singular patterns against noise, using the vector field shown in Fig. 9 (a), which contains six center points and two saddles. Here, to show the increasing noise, we draw these vector fields in the form of streamlines. As noise is added in the vector field (Fig. 9 (b)), our method can still identify all singularities without omission or false detection. In contrast, the triangular linear interpolation method [23] (Fig. 9(c)) fails to detect one center point and two saddle points while incorrectly identifying a spurious singularity (such a point has index 0 but is identified as a singularity). Moreover, the proposed method shows robustness in analyzing singular patterns in large-scale vector fields. As demonstrated in Fig. 10 (a), which shows streamlines of a noisy vector field comprising 10,000 vectors with multiple singularity types, our approach successfully identifies all singularities and correctly classifies their types, while in Fig. 10 (b), the triangular linear interpolation method does not find all singularities and identify their types correctly. These results clearly demonstrate the robustness of our approach for handling noise and large-scale vector fields.
Then we compare the mean positioning errors and running time between our method and several classical methods. To calculate the mean errors, suppose that we have data for $n$ moments, and we denote the $i$-th actual longitude as $al_{i}$ and the $i$-th positioning longitude as $pl_{i}$. The mean longitude error can be computed using the formula $\frac{1}{n}\sum_{i=1}^{n}\left|al_{i}-pl_{i}\right|$, and the mean latitude error can be computed similarly. Table 1 provides a comparison of the mean errors obtained using our method with other singularity extraction methods, they are methods based on triangular linear interpolation of vectors [23], computation of the Jacobian matrix [24], and Hodge decomposition [25] respectively. The average running times are also shown in Table 1. The Khanun datasets consists of 1600 sampling vectors (resulting in a digraph of 1600 vertices and 3120 edges), the Saola datasets comprises 1924 sampling vectors (resulting in a digraph of 1924 vertices and 3759 edges), and the geomagnetic field datasets consists of 1600 sampling vectors (resulting in a digraph of 1600 vertices and 3120 edges). Based on the results, our method achieves smaller errors than the method proposed by [25], and our time cost is lower than [25], which involves performing Hodge decomposition and locating local extrema of two energy functionals derived from the rotation-free and divergence-free components on the refined grids. Furthermore, our approach can achieve errors nearly as small as the numerical methods of [23] and [24], although it has a slightly longer time cost than those numerical methods. However, these numerical methods may fail to correctly identify singularities or encounter spurious singularities due to precision issues. For instance, in the experiments, both the triangular linear interpolation method and the Jacobian matrix method yield spurious singularities. The number of spurious singularities obtained from all vector fields in the corresponding datasets is shown in Table 1. For example, in the geomagnetic field datasets, there are 10 vector fields and 10 corresponding singularities in total, but the triangular interpolation method finds 15 singularities, including 5 spurious singularities, and the Jacobian method finds 12 singularities, including 2 spurious singularities. In contrast, our method, which relies on the singularity index, finds all of the singularities without any spurious ones. Additionally, our method has the capability to measure the variations between vector fields, while the other methods are unable to accomplish this.
Table 1: Comparative Results
Comparative Results for Khanun Datasets (36 vector fields)
Triangular interpolation [23] 0.134 0.107 27.21 15
Jacobian method [24] 0.126 0.109 45.25 9
Hodge decomposition [25] 0.205 0.147 566.64 0
our method 0.131 0.118 70.19 0
Comparative Results for Saola Datasets (32 vector fields)
Triangular interpolation [23] 0.108 0.093 34.76 15
Jacobian method [24] 0.111 0.093 56.00 2
Hodge decomposition [25] 0.129 0.133 689.39 0
our method 0.110 0.102 103.08 0
Triangular interpolation [23] 0.028 0.056 27.90 5
Jacobian method [24] 0.030 0.000 47.31 2
Hodge decomposition [25] 0.089 0.182 572.02 0
our method 0.037 0.059 90.95 0
6.5 Discussion
Our method exhibits several limitations. If the sampling grid is not sufficiently dense, some singularities may not be extracted because multiple singularities may exist within the same square, and their indices can cancel each other out. For example, if a singularity with index 1 and a singularity with index -1 locate in an identity square, the index of the square will finally becomes 0 and will not be identified as containing singularities. Furthermore, singularities located on the edges of the grid digraph may be overlooked, since the direction of this edge is unsure and the position of the singularity on the edge cannot be determined accurately. But in general, these situations are rarely to occur due to practical considerations and limitations in numerical accuracy. Additionally, since the generators of 2-dimensional path homology group is unclear currently, it is hard to generalize the proposed method to the vector fields in higher dimensional space.
7 Conclusions
In this study, we propose a novel method for determining the singularity within a planar vector field exhibiting a singular pattern. Moreover, we compare the changes of singular patterns of time-varying vector fields using persistence diagrams. Our approach involves the conversion of discrete vector fields into angle-based grid digraphs, which are then analyzed using persistent path homology. Experimental results demonstrate that our proposed method achieves low mean errors in locating the centers of tropical cyclones and locations of dip poles, and effectively measures the changes in tropical cyclone features over time.
In future work, we aim to explore alternative techniques for converting vector fields into digraphs to accommodate diverse applications. Furthermore, we will study how to address challenges related to vector fields with high-order singularities and vector fields in higher dimensional Euclidean spaces or surfaces, and investigate the potential utilization of higher dimensional (persistent) path homology as a new analytical tool.
Acknowledgments
CRediT authorship contribution statement
Yu Chen: Conceptualization, Investigation, Formal analysis, Methodology, Software, Validation, Writing – original draft, Writing – review & editing. Hongwei Lin: Conceptualization, Formal analysis, Methodology, Supervision, Funding acquisition, Writing – original draft, Writing – review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Appendices
Appendix A Proofs of some theorems and properties
A.1 Proof of the Theorem 3
Proof.
It is sufficient to prove that the singularities could not exist in the interior area of the boundary squares. Without loss of generality, let us consider the boundary square: ${C}$${D}$${A}$${B}$. According to the definition of the angle-based grid digraph, when moving from $A$ to $B$ and from $B$ to $D$, the vector undergoes a counterclockwise rotation, with the rotation angle ranging between $(0,\pi)$. Conversely, when moving from $D$ to $C$ and from $C$ to $A$, the vector rotates clockwise, with the rotation angle ranging between $(-\pi,0)$. By utilizing equation
$\operatorname{index}(L)=\frac{1}{2\pi}(\angle(\textbf{{v}}_{1},\textbf{{v}}_{2})+\angle(\textbf{{v}}_{2},\textbf{{v}}_{3})+\angle(\textbf{{v}}_{3},\textbf{{v}}_{4})+\angle(\textbf{{v}}_{4},\textbf{{v}}_{1}))$
we observe that along the closed curve $ABDCA$, the sum of rotation angles falls within the range of $(-2\pi,2\pi)$. Consequently, the number of revolutions made by the vector field $X$ while traveling around the curve $ABDCA$ is zero because zero is the only integer within the interval $(-1,1)$. As stated in Theorem 1 in main paper, this implies the absence of any singularities of $X$ within the interior of the boundary square. ∎
Remark 1.
In fact, for the square type ${C}$${D}$${A}$${B}$, its index is also zero since zero is the only integer within the interval $(-1,1)$. Thus this type of square will not contain singularity as well.
A.2 Proof of the Property 1 and Property 2
We first prove the following Lemma:
Lemma 1.
Proof.
A number of singular patterns such as sources, sinks, circulating and spiraling vector fields can be modeled or approximated by a field generated by logarithmic spirals [1]. The formula of each grid point can be discribed by
$\left\{\begin{matrix}x=ae^{\theta\cot\alpha}\cos\theta\\ y=\frac{a}{\rho}e^{\theta\cot\alpha}\sin\theta\end{matrix}\right.$
and the vector located on $(x,y)$ is
$\left\{\begin{matrix}P=ae^{\theta\cot\alpha}(\cot\alpha\cos\theta-\sin\theta)=x\cot\alpha-\rho y\\ Q=\frac{a}{\rho}e^{\theta\cot\alpha}(\cot\alpha\sin\theta+\cos\theta)=y\cot\alpha+\frac{x}{\rho}\end{matrix}\right.$
where $\theta$ is the polar angle, $\alpha\in[0,\pi)$ is the angle between the radial line and the tangent of the spiral at $(r,\theta)$; parameter $a$ is non-zero, and a counterclockwise field is generated by a positive $a$ while a clockwise field is generated by a negative $a$; parameter $\rho$ controls the width-to-height ratio.
To prove this theorem, we only need to consider the case of a horizontal line, as the vertical case can be proven similarly. Without loss of generality, let us assume that the singularity is located at $O(0,0)$. Suppose $(x,y)$ and $(x+\varepsilon,y)$ are two points in $U$, which can be described in polar coordinates as follows:
$x$ $=a_{1}e^{\theta_{1}\cot\alpha_{1}}\cos\theta_{1},\quad x+\varepsilon=b_{1}e^{\theta_{2}\cot\alpha_{2}}\cos\theta_{2}$
$y$ $=\frac{a_{1}}{\rho}e^{\theta_{1}\cot\alpha_{1}}\sin\theta_{1}=\frac{b_{1}}{\rho}e^{\theta_{2}\cot\alpha_{2}}\sin\theta_{2}$
Here, $a_{1}$, $b_{1}$, $\theta_{1}$, $\theta_{2}$, $\alpha_{1}$, $\alpha_{2}$ and $\rho$ are constants such that $y\neq 0$, with $a_{1}$ and $b_{1}$ both positive or negative. By rewriting the equation, we have:
$b_{1}e^{\theta_{2}\cot\alpha_{2}}$ $=b_{1}e^{\theta_{2}(\cot\alpha_{2}-\cot\alpha_{1}+\cot\alpha_{1})}$
$=(b_{1}e^{\theta_{2}(\cot\alpha_{2}-\cot\alpha_{1})})e^{\theta_{2}\cot\alpha_{1}}$
Let $a_{2}=b_{1}e^{\theta_{2}(\cot\alpha_{2}-\cot\alpha_{1})}$ and $\alpha=\alpha_{1}$. Then, we can express $x$, $x+\varepsilon$, $y$ as:
$x$ $=a_{1}e^{\theta_{1}\cot\alpha}\cos\theta_{1},\quad x+\varepsilon=a_{2}e^{\theta_{2}\cot\alpha}\cos\theta_{2}$
$y$ $=\frac{a_{1}}{\rho}e^{\theta_{1}\cot\alpha}\sin\theta_{1}=\frac{a_{2}}{\rho}e^{\theta_{2}\cot\alpha}\sin\theta_{2}$
Note that $a_{2}$ and $\alpha$ are also constants. Therefore, $\varepsilon$ can be written as:
$\varepsilon$ $=a_{2}e^{\theta_{2}\cot\alpha}\cos\theta_{2}-a_{1}e^{\theta_{1}\cot\alpha}\cos\theta_{1}$
$=a_{2}e^{\theta_{2}\cot\alpha}\cos\theta_{2}-a_{2}e^{\theta_{2}\cot\alpha}\frac{\sin\theta_{2}}{\sin\theta_{1}}\cos\theta_{1}$
$=a_{2}e^{\theta_{2}\cot\alpha}\frac{\sin(\theta_{1}-\theta_{2})}{\sin\theta_{1}}$
Now, let’s define:
$X_{1}$ $=(P(x,y),Q(x,y),0)$
$=(x\cot\alpha-\rho y,y\cot\alpha+\frac{x}{\rho},0)$
$X_{2}$ $=(P(x+\varepsilon,y),Q(x+\varepsilon,y),0)$
$=((x+\varepsilon)\cot\alpha-\rho y,y\cot\alpha+\frac{x+\varepsilon}{\rho},0)$
Thus, the third component of $X_{2}\times X_{1}$ is given by:
$z=$ $((x+\varepsilon)\cot\alpha-\rho y)(y\cot\alpha+\frac{x}{\rho})$
$-(x\cot\alpha-\rho y)(y\cot\alpha+\frac{x+\varepsilon}{\rho})$
$=$ $\varepsilon y(1+\cot^{2}\alpha)$
$=$ $a_{2}y(1+\cot^{2}\alpha)e^{\theta_{2}\cot\alpha}\frac{\sin(\theta_{1}-\theta_{2})}{\sin\theta_{1}}$
If $y>0$, then $\theta_{1}$ and $\theta_{2}$ lie in the interval $(0,\pi)$, and $\sin(\theta_{1}-\theta_{2})>0$ and $\sin\theta_{1}>0$ (see left part in Fig. 11), thus $z$ has the same sign as $a_{2}$. If $y<0$, then $\theta_{1}$ and $\theta_{2}$ lie in the interval $(\pi,2\pi)$, and $\sin(\theta_{1}-\theta_{2})<0$ and $\sin\theta_{1}<0$ (see right part in Fig. 11), thus $z$ has the opposite sign of $a_{2}$. Since the choice of $x$, $y$, and $\varepsilon$ is arbitrary, the sign of $z$ only depends on the sign of $y$ and $a_{2}$, leading to our conclusion. ∎
Fig. 11: (left) Situation of $y>0$; (right) Situation of $y<0$.
Now we prove the Property 1 and Property 2.
A.2.1 Proof of Property 1
Proof.
We can establish this property by considering the weights of the edges in a grid digraph, which correspond to the rotation angles between the vectors at their respective endpoints. Given that all edges in a grid digraph are of the same length, by Lemma 1, as we approach the singularity, the field vector on an edge undergoes a faster rotation. Consequently, the rotation angle between the vectors at both ends of an edge increases as the edge gets closer to the singularity. Therefore, the weight of an edge is greater when it is closer to the singularity. Hence, the edge with the locally greatest weight must be part of the smallest square containing the singularity. ∎
A.2.2 Proof of Property 2
Proof.
Before adding the edge with the greatest weight, according to Lemma 1, the representation containing the singularity must be either a or a configuration. Therefore, the edge with the greatest weight that is added last must be positioned in the middle of this hexagon. Regardless of the direction of the added edge, a new representation of the minimum generator of $H_{1}(G)$ emerges. Specifically, it is the smallest square that contains the singularity, taking the shape of . ∎
Appendix B Some tables of supplementary data
Table 2: Positioning Results of Magnetic Dip Poles and Errors